Use periodic table to prove that the total formula mass is conserved in this equation:
Na2CO3 + H2O = NaOH + H2CO3
1
Expert's answer
2020-09-29T06:49:02-0400

The balanced equation is:

Na2CO3 + 2H2O = 2NaOH + H2CO3

The mass of the reactants is given by:

((2 * 11) + 6 + (3 * 8)) + (2 * ((2 * 1) + 8)) = 72

The mass of the products is given by:

(2 * (11 + 8 + 1)) + ((2 * 1) + 6 + (3 * 8)) = 72

Given that the total mass of the reagents is equal to the mass of the product, it can be said that mass has neither been lost nor gained in the reaction.